# Attainability of quantum state discrimination bounds with collective measurements on finite copies

### v1

@@ V Conclusion
We have obtained analytic expressions for the Helstrom bound for several simple qubit examples. Based on this we have evaluated the attainability of the asymptotic limits in quantum metrology. Our results indicate that for certain classes of qubit state discrimination problems the quantum Chernoff bound is approached only very slowly. In particular, in theorem 1, we have shown that, for certain qubit problems, the quantum Chernoff bound error exponent cannot be exactly saturated for any finite $M$. While this result is not particularly surprising, and is consistent with existing results pereira2023analytical ; audenaert2012quantum ; rouze2017finite ; li2014second , it is worthwhile to verify.
We have also investigated the conditions necessary to saturate the Helstrom bound. In particular we have investigated a family of measurements which generate entanglement on less than all $M$ copies of the unknown state. We did not find any instances where this class of measurement saturates the Helstrom bound. However, there does appear to exist a hierarchy within this family of measurements. Additionally, we have obtained initial results suggesting that LOCC measurements are more powerful than repeated 2-copy entangling measurements without classical communication. However whether this is a general result or not requires further investigation.
Finally, it could be useful to extend our results to other settings, particularly the discrimination of Gaussian states lloyd2008enhanced ; tan2008quantum ; bradshaw2021optimal , or the use of other figures of merit, such as asymmetric state discrimination karsa2020quantum ; pereira2023analytical or unambiguous state discrimination ivanovic1987differentiate ; dieks1988overlap . Given the strong connection between quantum classification and state discrimination banchi2021generalization , it would be interesting to evaluate the performance of the family of measurements introduced here to this and related fields, such as the learning of quantum processes seif2024entanglement .
@@ VI Methods
The parametrized quantum circuits comprise alternating layers of $U(2)$ unitary gates on every qubit and CNOT gates between pairs of adjacent qubits. The CNOTs are in brickwork pattern with closed (i.e. wrap-around) boundary conditions at the ends of the qubit chain, such that on an $N$-qubit circuit, a minimum of $\sim N/2$ CNOT layers suffice to generate an entanglement light cone spanning all qubits. Each $U(2)$ gate is parametrized by $3$ angles. We minimized the error probability (see Eqs. (38) and (42) of main text) over all $U(2)$ gate angles, using a standard optimization algorithm (L-BFGS-B) with basin hopping. Our calculations employed 50 hops with a maximum of 500 iterations each. We checked for convergence by increasing the number of CNOT layers in the parametrized quantum circuits (up to  14), and number of optimization hops/iterations, to ensure that the best-found error probability is stable.
@@ Acknowledgements
For simultaneously diagonalisable states the probability of error is given by audenaert2014upper
@@ Appendix B Necessary and sufficient condition for a POVM to be saturate the Helstrom bound
Following Ref. ban1997optimum , we note that for distinguishing any two states $\rho_{1}$ and $\rho_{2}$ we can use a two outcome POVM, $\{\Pi_{1},\Pi_{2}\}$. In order to saturate the Helstrom bound, this POVM must satisfy
@@
Appendix C Comparison of the error exponents as a function of number of copies
For reference we also plot $R_{\epsilon}$ here, shown in Fig. 9.
Figure 9: Scaling of error exponents for a simple qubit example. a) We plot the the normalised ratio of the $M$-copy exponent to the asymptotic limit. The solid lines correspond to even $M$ and the dashed lines to odd $M$. b) We plot the difference between the $M$ and $M+2$-copy exponents. The trend is shown only for even $M$, but the trend for odd $M$ is similar.
Appendix D Calculating the error probability when repeatedly performing the $M$-copy Helstrom measurement
We now analyse the measurement strategy discussed in section III.1.5, where we have $N$ copies of the quantum state and we perform the $M$-copy Helstrom measurement $N/M$ times ($N\gg M$). Eq. (19) gives an analytic expression for the error probability when the Helstrom measurement is implemented. From the symmetry of the problem being considered (equal prior probabilities) we can conclude that the Helstrom measurement is symmetric. Hence, the probability of each outcome of the Helstrom measurement is equal to $P^{\text{H}}_{\text{e}}(M)$. We then repeat this measurement $N/M$ times, which can be modelled as a binomial distribution. Let us consider the following strategy. We assign the value 1 to successful events and the value 0 to incorrectly identified density matrices. In this case the expected value from the sum of all the random variables ($X=\sum_{i=1}^{N/M}X_{i}$) is $\mu=(1-P^{\text{H}}_{\text{e}}(M))N/M>N/2M$, and we will incorrectly guess the unknown density matrix if the sum of the random variables has a value less than $N/2M$. The classical Chernoff bound for the sum of $n$ repeated binomially distributed random variables characterised by probability $p$ states that arratia1989tutorial
@@ Appendix D Calculating the error probability when repeatedly performing the $M$-copy Helstrom measurement
Appendix E Unattainability of the Chernoff bound for pure qubit states
Denote $\gamma=\text{cos}(\alpha)^{2M}$,
For $10<\gamma<$ 1, $\sqrt{1-\gamma}>1-\gamma$.
Appendix F LOCC measurement saturating the Helstrom bound for two copies of the unknown state and equal prior probabilities

### v2

@@ V Conclusion
We have obtained analytic expressions for the Helstrom bound and Chernoff bound for several simple qubit examples. Based on this we have evaluated the attainability of the asymptotic limits in quantum state discrimination. In particular we examined whether repeatedly implementing the $M$-copy Helstrom measurement is sufficient to approach the Chernoff bound asymptotically. In this setting we found that the single-copy Helstrom measurement performed best. However, this comes with the caveat that for repeatedly implementing the same measurement other measurements will be optimal and one could expect that some other repeated entangling measurement will outperform the single-copy measurement. However, finding this optimal measurement is left as an open quesiton.
We have also investigated the conditions necessary to saturate the Helstrom bound. In particular we have investigated a family of LOCC measurements which generate entanglement on less than all $M$ copies of the unknown state. We did not find any instances where this class of measurement saturates the Helstrom bound. However, there does appear to exist a hierarchy within this family of measurements. Similar results were obtained for a family of PPT measurements. However, as is shown in Fig. 6, the attainability of the Helstrom bound depends strongly on the structure of the problem. Further investigation is needed to find the conditions under which the Helstrom bound is attainable without entanglement across all available modes.
Finally, it could be useful to extend our results to other settings, particularly the discrimination of Gaussian states [69, 70, 71], or the use of other figures of merit, such as asymmetric state discrimination [72, 45] or unambiguous state discrimination [73, 74]. Given the strong connection between quantum classification and state discrimination [75], it would be interesting to evaluate the performance of the family of measurements introduced here to this and related fields, such as the learning of quantum processes [76].
@@ VI Methods
The parametrized quantum circuits comprise alternating layers of $U(2)$ unitary gates on every qubit and CNOT gates between pairs of adjacent qubits, shown in Fig. 7. The CNOTs are in brickwork pattern with closed (i.e. wrap-around) boundary conditions at the ends of the qubit chain, such that on an $N$-qubit circuit, a minimum of $\sim N/2$ CNOT layers suffice to generate an entanglement light cone spanning all qubits. Each $U(2)$ gate is parametrized by $3$ angles. We minimized the error probability (see Eqs. (20) and (24) of main text) over all $U(2)$ gate angles, using a standard optimization algorithm (L-BFGS-B) with basin hopping. Our calculations employed 50 hops with a maximum of 500 iterations each. We checked for convergence by increasing the number of CNOT layers in the parametrized quantum circuits (up to  14), and number of optimization hops/iterations, to ensure that the best-found error probability is stable.
Figure 7: Circuit structure used to find optimal LOCC measurements in section IV.3. An example of the circuit structure used for 6 qubits with 3 layers is shown. Blue boxes represent single qubit operations and purple boxes represent CNOT gates.
@@ Acknowledgements
For simultaneously diagonalisable states the probability of error is given by [36]
@@ Appendix B Necessary and sufficient condition for a POVM to be saturate the Helstrom bound
Following Ref. [37], we note that for distinguishing any two states $\rho_{1}$ and $\rho_{2}$ we can use a two outcome POVM, $\{\Pi_{1},\Pi_{2}\}$. In order to saturate the Helstrom bound, this POVM must satisfy
@@
Appendix C Calculating the error probability when repeatedly performing the $M$-copy Helstrom measurement
We now analyse the measurement strategy discussed in section III.1.4, where we have $N$ copies of the quantum state and we perform the $M$-copy Helstrom measurement $N/M$ times ($N\gg M$). Eq. (9) gives an analytic expression for the error probability when the Helstrom measurement is implemented. From the symmetry of the problem being considered (equal prior probabilities) we can conclude that the Helstrom measurement is symmetric. Hence, the probability of each outcome of the Helstrom measurement is equal to $P^{\text{H}}_{\text{e}}(M)$. We then repeat this measurement $N/M$ times, which can be modelled as a binomial distribution. Let us consider the following strategy. We assign the value 1 to successful events and the value 0 to incorrectly identified density matrices. In this case the expected value from the sum of all the random variables ($X=\sum_{i=1}^{N/M}X_{i}$) is $\mu=(1-P^{\text{H}}_{\text{e}}(M))N/M>N/2M$, and we will incorrectly guess the unknown density matrix if the sum of the random variables has a value less than $N/2M$. The classical Chernoff bound for the sum of $n$ repeated binomially distributed random variables characterised by probability $p$ states that [77]
@@ Appendix C Calculating the error probability when repeatedly performing the $M$-copy Helstrom measurement
Appendix D LOCC measurement saturating the Helstrom bound for two copies of the unknown state and equal prior probabilities
@@
Appendix E Quantities to investigate the error scaling in finite limit
Finally, we wish to note that in the initial stages of this study we considered alternative methods of quantifying the attainability of the quantum Chernoff bound, which we present here for any interested reader. Essentially, we have investigated whether or not the quantum Chernoff bound is attainable given any finite number of copies of the unknown state. This amounts to asking whether the error probability provided by the Helstrom bound (when appropriately normalised for the number of copies of the unknown state being considered) ever reaches the error probability specified by the quantum Chernoff bound. However, the physical interpretation of these results is not so obvious as the Chernoff bound is exclusively an asymptotic quantity.
Our question is whether the error probability given by the Chernoff bound is attainable with any finite number of copies of the unknown state. To clarify what we mean by this, we have plotted two functions in Fig. 8. For one of these functions, the asymptotic limit of the function is attainable for finite arguments, for the other it is not. As the Chernoff bound specifically applies only in the asymptotic limit, it may be expected that the asymptotic limit cannot be reached for any finite argument, as is the case for many mathematical functions. Nevertheless, it is still worthwhile to investigate this, and related, questions. For instance, when investigating this question, we found interesting results about the rate at which the Chernoff bound is approached.
In order to address this question we will need some method of normalising between the Helstrom bound (which applies in the finite limit) and the Chernoff bound (which applies in the asymptotic limit). Noting that
$\lim_{M\to\infty}P_{\text{e,min},M}^{1/M}=\kappa\;.$ (40)
Note that for distinguishing more than two states, the Chernoff bound can be off by a constant factor [36]. However, for distinguishing two states this effect is not present. By comparing $\epsilon_{M}$ and $\epsilon_{\infty}$ we can examine how rapidly the ultimate limits on the error exponent can be approached.
When distinguishing quantum states we find situations where $\epsilon_{1}\geq\epsilon_{2}\geq\epsilon_{3}\ldots\geq\epsilon_{\infty}$, or equivalently $P_{\text{e}}(1)\leq P_{\text{e}}(2)^{1/2}\leq P_{\text{e}}(3)^{1/3}\ldots\leq\kappa$. At first glance this may appear surprising, however this is entirely in keeping with the results of Refs. [40, 41]. Specifically Ref. [41] proved
and Ref. [40] derived
which implies
$\frac{\text{log}(P_{\text{e,min},M})}{M}\leq\text{log}(\kappa)\;.$ (44)
Let us define
$\delta(M,N)=R(M)-R(N)\;,$ (46)
E.1 Finite size comparison of bounds in example 1
Given Eqs. (9) and (13), we are in a position to evaluate the quantities $R(M)$ and $\delta(M,N)$. In Fig. 9 a), we plot $1-R(M)$ as a function of $M$. Although it is not possible to say definitively that $R(M)$ does not plateau to 1 for any finite $M$, it certainly appears that way from this figure. In Fig. 9 b), we plot $\delta(M+2,M)$ as a function of $M$. In this case it is evident that this function is continuously decreasing. However, again an analytic proof of this statement is not presented here.
As this example is essentially classical (the two states are simultaneously diagonalisable), this appears to suggest that the inability to exactly saturate the Chernoff bound is not a quantum property. Indeed similar concepts have been studied in classical information theory, where second order corrections to the asymptotic limits are known [78].
E.2 Finite sample comparison of bounds in example 2
Using Eqs. (18) and  (15), we can evaluate the normalised ratio of the Helstrom bound to the Chernoff bound (Eq. (45)) as
Below we prove that this ratio never reaches the value 1 for any finite $M$, hence the Chernoff bound cannot be exactly saturated. This is consistent with existing results for asymmetric quantum state discrimination [45, 46, 47, 48]. Of course it is still possible that the Chernoff bound can be approximately saturated, as is shown in Fig. 10. In this figure, we plot 1 minus the normalised ratio of the $M$-copy error probability to the asymptotic error probability. This plot shows that the ratio is an increasing function of $M$, however the ratio never reaches exactly 1. Therefore, the asymptotic limit cannot be reached with any finite-copy measurement. This result is stated as the following theorem
Theorem 1
Proof.
Denote $\gamma=\text{cos}(\alpha)^{2M}$,
For $10<\gamma<$ 1, $\sqrt{1-\gamma}>1-\gamma$.
E.2.1 Numerical simulations for mixed states